OpenStudy (lmvillamagua):

Hey guys, Can you help me solving the following equation? cosx - sin2x =0 Thanks, Luis

OpenStudy (the_puppet1):

Uh what?

OpenStudy (griefer101):

wow thats complicated

OpenStudy (thenaughtyserperior):

2sinxcosx - cosx = 0 cosx(2sinx - 1) = 0 cosx = 0 or sinx = 1/2 x = pi/2, 3pi/2, pi/2+2kpi, 3pi/2+2kpi x= pi/6, 5pi/6, pi/6+2kpi, 5pi/6+2kpi where k is any integer

OpenStudy (griefer101):

graphing calculator

OpenStudy (griefer101):

my brain just died